# Investigation of the effects of superhydrophobic surface treatment on the dynamics of the flow in the near wake of a sphere using spatial dynamic mode decomposition

### v1

@@ Abstract
Viscous drag results from the fluid at a surface having zero relative velocity, a phenomenon known as the no-slip condition. Superhydrophobic surfaces, when submerged in water, trap a layer of air in their surface texture, partially replacing the liquid-solid interface with a liquid-gas interface. This air layer, called the plastron, results in partial slip at the surface, thereby reducing the viscous drag. In turbulent flows, large fluctuations in pressure and velocity can deplete or completely remove the plastron from the surface. This makes evaluating the effects of superhydrophobic surface treatments on flow dynamics particularly challenging. This study examines the impact of a sustained plastron on flow dynamics in the wake of a sphere, achieved by supplying air at low pressure through pores in the sphere’s surface. Instantaneous planar velocities in the wakes of spheres, both with and without superhydrophobic surface treatment, are measured within a plane passing through the sphere’s centre. Dynamic mode decomposition is applied to the velocity fluctuations in the wake to evaluate how superhydrophobic surface treatments affect flow dynamics in the near wake. It is shown that the addition of the pores has a relatively small effect on the dynamics in the wake of the sphere, while the dynamics in the wake are radically changed by the addition of superhydrophobic surface treatment when the plastron is sustained.
@@ 1.1 Superhydrophobic Surfaces
The lotus leaf exhibits a unique ability to repel water due to a combination of surface roughness and chemical properties, enabling water droplets to roll off without wetting the surface [1, 2]. This phenomenon, termed superhydrophobicity, has garnered significant interest for its potential applications, particularly in drag reduction and anti-fouling technologies [3, 4]. Water droplets on a superhydrophobic surface form with a large contact angle, which is the internal angle between the surface and the edge of the droplet [5, 6]. When sliding down an inclined surface, the droplets exhibit low contact angle hysteresis, which is the difference between the angle of the droplet at its leading and trailing edges [7, 3]. The texture of superhydrophobic surfaces consists of multi-scale roughness, which amplifies the effect of chemical hydrophobicity [8]. Due to this combination of properties, the droplets are easily perturbed [9, 10, 11, 12].
@@
These properties are achieved by trapping a layer of air, known as the plastron, between the surface and the droplet in what is known as the Cassie-Baxter state [13]. When submerged in water, the air trapped at the surface leads to a partial air-water interface and thus a partial slip over the surface [3], which reduces the skin friction at the surface and decreases the viscous drag of flow over the surface. A major factor in the varying results of experiments with superhydrophobic surfaces in turbulent flows is the propensity for the plastron to be depleted or completely deteriorated by the large velocity and pressure fluctuations present in turbulent flows. Even in cases where the plastron remains attached to the surface, these fluctuations have a significant effect on its distribution over the surface [14]. Once the plastron is depleted, the surface roughness begins to protrude into the flow, and the drag increases. This can occur across the entire surface as the plastron deteriorates, or in local regions as the plastron is deformed by the forces exerted on it by the flow.
@@
As a plastron-covered superhydrophobic surface decreases viscous drag in water flows, it is of particular interest in reducing fuel consumption in marine vessels, such as ships and submarines, and for reducing transport costs in water supply pipes. While the drag reduction of superhydrophobic surfaces in laminar flows is directly related to the amount of slip produced, which is determined by the fraction of the surface aerated by the plastron [3], drag reduction in turbulent flows is more complicated to determine. While drag reduction has been achieved in turbulent flows, such as boundary layers [15] and Taylor-Couette flows [16], investigations of superhydrophobic drag reduction have also resulted in negligible changes [17], or even significant increases [18] in drag. The depletion of the plastron in turbulent boundary layers at high Reynolds number in the study by Aljallis et al. [19] resulted in an increase in drag, while the reduction of drag on a boat in open water observed by Xu et al. [20] relied on the plastron being sustained. The study of turbulent boundary layers on superhydrophobic surfaces by Gose et al. [21], over a Reynolds number range of 10,000 to 30,000, resulted in drag reduction of 90% to drag increases of 90%. This depended on the Reynolds number and surface properties, with the product of the surface-to-projected area ratio of the wetted region and the proportion of the surface area that is aerated being the best predictor of drag reduction.
@@ 1.2 Flow Structures in the Wake of A Sphere
As the application of superhydrophobic surfaces for drag reduction is of particular interest for maritime vessels, such as ships and submarines, the effects of superhydrophobic surface treatment on the flow over bluff bodies are an important consideration to assess the efficacy of these surfaces. Spheres serve as an idealised model for axisymmetric bluff bodies, exhibiting reverse flow, flow separation, and periodic and oscillatory vortex shedding in their wakes [22, 23, 24]. Quantitative and qualitative measurements in the wake of the sphere have identified numerous structures and characterised their development over various Reynolds numbers. The wake loses axisymmetry around $Re_{D}\approx 20$, with a vortex ring forming at $Re_{D}\approx 25$ [22, 25]. The vortex ring begins to oscillate at $Re_{D}\approx 130$, and elongates as the Reynolds number increases. Single-frequency vortex shedding begins at $Re_{D}\approx 300$ [26, 27, 24], and the generation of a lateral force [28, 29] results in the loss of planar symmetry by $Re_{D}\approx 500$. Small-scale vortical structures are present within the main vortex ring from $Re_{D}\approx 1,000$ as the vortex rings are shed from the sphere and break down as they move downstream [29, 30].
@@ 1.3 The Current Work
The current work investigates the effect of superhydrophobic surface treatment on the dynamics of the flow in the near wake of a sphere, with the plastron actively maintained by supplying air through pores in the surface of the sphere. Three spheres were manufactured: (i) a smooth untreated sphere to serve as a reference, (ii) an untreated sphere with a porous shell to determine the change in dynamics from the addition of the pores, and (iii) a porous sphere with superhydrophobic surface treatment to determine the change in the flow dynamics when the plastron is sustained.
The instantaneous velocities in a plane through the near wake of each sphere were measured using particle image velocimetry (PIV), from which the instantaneous velocity fluctuations were determined. Spatial dynamic mode decomposition (DMD) was performed in the streamwise direction on the velocity fluctuations in the near wake of each sphere. The resulting DMD modes were sorted from slowest to fastest decay rate. The slowest decaying DMD modes for each sphere were compared to determine the effect of the superhydrophobic surface treatment on the dynamics of the near wake when the plastron is maintained.
The results show that the addition of pores to the sphere has relatively little effect on the slowest-decaying DMD modes present in the near wake, although there are changes to the decay rates and wavelengths of these modes. In contrast, the addition of superhydrophobic surface treatment radically changes the DMD modes in the wake due to the distribution of the plastron over the sphere. While the three slowest-decaying DMD modes in the wakes of the untreated spheres exhibit planar symmetry, only the slowest-decaying DMD mode of the sphere with superhydrophobic surface treatment is planar symmetric. The effects of the increase in three-dimensionality in the flow on the plastron are likely to be highly dependent on the flow geometry and Reynolds number, and will affect the efficacy of the superhydrophobic surface treatment at reducing drag. Varying geometries and flow conditions are also likely to have a substantial effect on how the addition of superhydrophobic surface treatments affects the dynamics of the flow, which may enhance or mitigate the drag reduction effects of the partial slip introduced by the superhydrophobic surface.
@@ 2.1.1 Vertical Water Tunnel
The flow in the near wake of the spheres was measured in a vertical water tunnel facility with a 250 mm $\times$ 250 mm cross-section and a 1.5 m working section [36, 37, 38], as shown in figure 1a. The settling chamber above the working section includes a 16:1 contraction and honeycomb screens to reduce the turbulence in the working section. The 0.5 m high test section is located in the middle of the working section, with a cut-out section on one wall to provide access. An acrylic panel is used to seal the tunnel, and is flush with the internal side of the wall to minimise the disturbance to the flow. Water is pumped from the plenum chamber below the working section up to the settling chamber through a return pipe, such that the flow through the working section is downwards. The introduction of seeding to the flow and the filtration of the tunnel are performed using an auxiliary circuit connected to the return pipe, which is isolated from the tunnel during experiments.
@@
Figure 1: (a) Vertical water tunnel facility [38] and (b) experimental field of view.
Figure 2: Aerofoil crossbeam and sting used for mounting the sphere [38].
@@ 2.1.2 Sphere Design
Figure 3: Cut-away view of 3D-printed sphere models for the (a) smooth reference sphere and (c) porous spheres with and without superhydrophobic surface treatment [38].
@@ 2.2 Experimental Method
The freestream velocity of the tunnel was set to $U_{\infty}=200$ mm/s for all spheres, resulting in a Reynolds number of $Re_{D}=7,780$, with a corresponding background turbulence of 2% [38]. The mean convective time-scale arising from the sphere diameter and freestream velocity is $t_{c}=0.2$[38]. The flow was seeded with glass spheres of 11µm diameter, with the single-exposed particle images illuminated by a pair of 400 mJ Spectra Physics Nd:YAG lasers firing 7 ns pulses and recorded using a PCO Panda 26 DS camera. The camera consists of a 5120 $\times$ 5120 array of 2.5 µm $\times$ 2.5 µm pixels, which was used to capture a 190 mm $\times$ 190 mm field of view beginning directly below the sphere and aligned to its centre, as shown in figure 1b, with a resulting spatial resolution of 37 µm/px. Pairs of single-exposed images were captured with a separation of $\Delta t=1$ ms, and image pairs were acquired at a rate of 1 Hz, resulting in a separation between instantaneous velocity fields of $5t_{c}$. The acquisition of $N_{T}>10,000$ images for each sphere was synchronised with the laser pulses using a BeagleBone Black (BBB) pulse generator [39].
The plastron formed on the SHS sphere was maintained by supplying air at low pressure, with the pores cleared of any settled seeding particles periodically [38]. The flow rate required to maintain the plastron without causing air to jet from the pores was 0.18$\pm$0.006 litres per minute. No roughness was added to the sphere, and as a result the plastron was thinner than those in settling sphere experiments with added roughness [14], as shown in figure 4. This reduced the risk of reflections into the camera in the field of view.
Figure 4: Image of the plastron formed on the superhydrophobic sphere, taken with a non-PIV camera [38].
@@
Freestream velocity $U_{\infty}$ 200 mm/s --
Sphere diameter $D$ 40 mm --
Reynolds number $Re_{D}$ 7,780 --
@@
[…]
Number of 2C-2D velocity fields $N_{T}$ $>$10,000
The uncertainty of the 2C-2D-MCCDPIV algorithm used is 0.06 px [40], which corresponds to a single-sample uncertainty of 2.22 mm/s in the instantaneous velocity $u$, due to random error. The mean of 10,000 instantaneous velocity vectors, $\overline{u}$, therefore has an uncertainty of 0.0222 mm/s. As the instantaneous velocity fluctuations are given by
@@
and is equal to 1.1% of the freestream velocity.
@@ 2.3 Dynamic Mode Decomposition
Modal decomposition is commonly used to determine the time information in patterns in turbulent fluctuations, which may indicate coherent structures and their time-scale in the flow. The basis of modal decomposition is the singular value decomposition (SVD), which is also the basis of proper orthogonal decomposition (POD[41, 42]. The SVD of the $m\times n$ matrix $\boldsymbol{X}$, where $\boldsymbol{X}_{m,n}$ represents a scalar field at time $t_{m}$ and spatial coordinate $\boldsymbol{x}_{n}$, is given by
@@
The matrix $\boldsymbol{X}_{x}$ was defined using the instantaneous velocity fluctuations at all streamwise locations except for the furthest downstream, where each row represents an instant in time and each column represents a position in space. In order for both velocity components to be included simultaneously, the matrix $\boldsymbol{X}_{x}$ was defined as
@@
where $t_{m}$ is the time of the $m^{th}$ velocity field, ${(x,y)}_{n}$ is the $n^{th}$ spatial coordinate, and $N$ denotes the total number of spatial coordinates excluding those at the furthest downstream location. The matrix $\boldsymbol{X}_{x+\Delta x}$ was thus defined by replacing $x$ with $x+\Delta x$ in equation (10).
The SVD of $\boldsymbol{X}_{x}$ was computed using equation (6) and the reduced-rank SVD was defined as
@@
where $r$ is the rank of the reduction, and the rank-reduction ratio is given by
$RRR=\frac{\Sigma_{r}^{2}}{\Sigma_{1}^{2}}.$ (12)
A rank-reduction ratio of $RRR=0.25$ was chosen for all spheres to limit the DMD to the most dominant POD modes. This resulted in ranks of ten, nine, and seven for the REF, PRS, and SHS spheres, respectively. The cumulative and relative turbulent kinetic energy of the included POD modes are shown in figure 5 [38].
Figure 5: (a) Cumulative turbulent kinetic energy of the leading POD modes and (b) the ratio of the energy of the leading POD modes to the first POD mode for each sphere [38]. Included modes are shown by solid shapes and excluded modes are shown by hollow shapes.
The reduced-rank SVD of $\boldsymbol{X}_{x}$ and the full matrix $\boldsymbol{X}_{x+\Delta x}$ were then used to determine the matrix $\boldsymbol{\widetilde{A}}$ using equation (8). The eigendecomposition of $\boldsymbol{\widetilde{A}}$
@@
yields eigenvalues $\widetilde{\mu}$ and eigenvectors $\boldsymbol{\Phi}$, from which the DMD modes of $\boldsymbol{\widetilde{A}}$ are given by
@@
where $k$ is an integer from 1 to $r$. The eigenvalues of $\boldsymbol{A}$, $\mu$, are related to those of $\boldsymbol{\widetilde{A}}$ by
@@
where $|\widetilde{\mu}_{k}|$ and $\theta_{\widetilde{\mu}_{k}}$ are the magnitude and angle of $\widetilde{\mu}_{k}$ in the complex plane, respectively. The wavenumber defined in equation (17) is in radians per metre and is divided by $2\pi$ to be converted to m-1, therefore the wavelength of the $k^{th}$ DMD mode in the streamwise direction, $\lambda^{x}_{k}$ is
@@
The eigenvalues of $\boldsymbol{A}$ and $\boldsymbol{\widetilde{A}}$ are shown in figures 6a and 6b, respectively. The decay rates and wavelengths of the three slowest-decaying DMD modes are given in table 2, and the corresponding DMD modes in the wake of each sphere are shown in figures 78, and 9, respectively. In order to provide an additional visualisation of the structures in the DMD modes, the out-of-plane vorticity
$\omega^{\prime}(x,y)=\frac{\partial v^{\prime}(x,y)}{\partial x}-\frac{\partial u^{\prime}(x,y)}{\partial y},$ (19)
of the real and imaginary parts of each DMD mode was calculated. The resulting vorticity fields in the wake of each sphere are shown in figures 1011, and 12, for the first, second, and third DMD modes of each sphere, respectively.

### v2

@@ Abstract
Viscous drag arises from the fluid at a surface having zero relative velocity, a phenomenon known as the no-slip condition. Superhydrophobic surfaces, when submerged in water, trap a layer of air in their surface texture, partially replacing the liquid-solid interface with a liquid-gas interface. This air layer, called the plastron, results in partial slip at the surface, thereby reducing the viscous drag. In turbulent flows, large fluctuations in pressure and velocity can deplete or completely remove the plastron from the surface. This makes evaluating the effects of superhydrophobic surface treatments on flow dynamics particularly challenging. This study examines the impact of a sustained plastron on the dynamics in the shear layer of a sphere, achieved by supplying air at low pressure through pores in the sphere’s surface. Instantaneous planar velocities in the wakes of spheres, both with and without superhydrophobic surface treatment, are measured within a plane passing through the spheres’ centre. Dynamic mode decomposition (DMD) is applied to the velocity fluctuations in the shear layer to evaluate how superhydrophobic surface treatment affects the instabilities there. It is shown that the addition of the pores has a relatively small effect on the instabilities in the shear layer, while they are significantly changed by the addition of superhydrophobic surface treatment when the plastron is sustained.
Keywords: flow over spheres, wakes, superhydrophobic surfaces, dynamic mode decomposition
@@ 1.1 Superhydrophobic Surfaces
The lotus leaf exhibits a unique ability to repel water due to a combination of surface roughness and chemical properties, enabling water droplets to roll off without wetting the surface [1, 2]. This phenomenon, termed superhydrophobicity, has garnered significant interest for its potential applications, particularly in drag reduction and anti-fouling technologies [3, 4]. Water droplets on a superhydrophobic surface form with a large contact angle, which is the internal angle between the surface and the edge of the droplet [5, 6]. When sliding down an inclined surface, the droplets exhibit low contact angle hysteresis, which is the difference between the angle of the droplet at its leading and trailing edges [7, 3]. The texture of superhydrophobic surfaces consists of multiscale roughness, which amplifies the effect of chemical hydrophobicity [8]. Due to this combination of properties, the droplets are easily perturbed [9, 10, 11, 12].
@@
These properties are achieved by trapping a layer of air, known as the plastron, between the surface and the droplet in what is known as the Cassie-Baxter state [13]. When submerged in water, the air trapped at the surface leads to a partial air-water interface and thus a partial slip over the surface [3], which reduces the skin friction at the surface and decreases the viscous drag of flow over the surface. A major factor in the varying results of experiments with superhydrophobic surfaces in turbulent flows is the propensity for the plastron to be depleted or completely deteriorated by the large velocity and pressure fluctuations present in turbulent flows. Even in cases where the plastron remains attached to the surface, these fluctuations have a significant effect on its distribution over the surface [14]. This was shown to affect the transverse velocity of a settling sphere by Castagna et al. [15]. Once the plastron is depleted, the surface roughness begins to protrude into the flow, and the drag increases. This can occur across the entire surface as the plastron deteriorates, or in local regions as the plastron is deformed by the forces exerted on it by the flow.
@@
As a plastron-covered superhydrophobic surface decreases viscous drag in water flows, it is of particular interest in reducing fuel consumption in marine vessels, such as ships and submarines, and for reducing transport costs in water supply pipes. While the drag reduction of superhydrophobic surfaces in laminar flows is directly related to the amount of slip produced, which is determined by the fraction of the surface aerated by the plastron [3], drag reduction in turbulent flows is more complicated to determine. While drag reduction has been achieved in turbulent flows, such as boundary layers [16] and Taylor-Couette flows [17], investigations of superhydrophobic drag reduction have also resulted in negligible changes [18], or even significant increases [19] in drag. The depletion of the plastron in turbulent boundary layers at high Reynolds number in the study by Aljallis et al. [20] resulted in an increase in drag, while the reduction of drag on a boat in open water observed by Xu et al. [21] relied on the plastron being sustained. The study of turbulent boundary layers on superhydrophobic surfaces by Gose et al. [22], over a Reynolds number range of 10,000 to 30,000, resulted in drag effects ranging from a 90% reduction to a 90% increase. This depended on the Reynolds number and surface properties, with the product of the surface-to-projected area ratio of the wetted region and the proportion of the surface area that is aerated being the best predictor of drag reduction.
@@ 1.2 Flow Structures in the Wake of A Sphere
As the application of superhydrophobic surfaces for drag reduction is of particular interest for maritime vessels, such as ships and submarines, the effects of superhydrophobic surface treatment on the flow over bluff bodies are an important consideration to assess the efficacy of these surfaces. Spheres serve as an idealised model for axisymmetric bluff bodies, exhibiting reverse flow, flow separation, and periodic and oscillatory vortex shedding in their wakes [23, 24, 25]. Quantitative and qualitative measurements in the wake of the sphere have identified numerous structures and characterised their development over various Reynolds numbers. The wake loses axisymmetry around $Re_{D}\approx 20$, with a vortex ring forming at $Re_{D}\approx 25$ [23, 26]. The vortex ring begins to oscillate at $Re_{D}\approx 130$, and elongates as the Reynolds number increases. Axisymmetry in the wake is broken by a pitchfork bifurcation around $Re_{D}=212$ [27]. A Hopf bifurcation is formed around $Re_{D}=275$ [27] after which the flow is dominated by single-frequency vortex shedding [28, 29, 25]. A quasiperiodic regime with a second dominant frequency begins at $350<Re_{D}<450$, after the emergence of a Neimark-Sacker bifurcation around $Re_{D}=330$ [27]. The generation of a lateral force [30, 31] results in the loss of planar symmetry by $Re_{D}\approx 500$. Small-scale vortical structures are present within the main vortex ring from $Re_{D}\approx 1,000$ as the vortex rings are shed from the sphere and break down as they move downstream [31, 32].
@@ 1.3 The Current Work
The current work investigates the effect of superhydrophobic surface treatment on the dynamics of the shear layer in the near wake of a sphere, where the plastron is actively sustained by supplying air through pores in the surface of the sphere. Three spheres were manufactured: (i) a smooth untreated sphere to serve as a reference, (ii) an untreated sphere with a porous shell to determine the change in dynamics from the addition of the pores, and (iii) a porous sphere with superhydrophobic surface treatment to determine the change in the flow dynamics when the plastron is sustained. The instantaneous velocities in a plane aligned with the centre of each sphere were measured using particle image velocimetry (PIV), from which the instantaneous velocity fluctuations were determined. Spatial dynamic mode decomposition (DMD) was performed in the streamwise direction on the velocity fluctuations in the shear layer of each sphere. The resulting DMD modes were sorted in order of descending energy share. The DMD modes for each sphere were compared to determine the effect of the superhydrophobic surface treatment on the dynamics when the plastron is sustained.
@@ 2.1.1 Vertical Water Tunnel
The flow in the near wake of the spheres was measured in a vertical water tunnel facility with a 250 mm $\times$ 250 mm cross-section and a 1.5 m working section [38, 39, 40], as shown in figure 1a. The settling chamber above the working section includes a 16:1 contraction and honeycomb screens to reduce the turbulence in the working section. The 0.5 m high test section is located in the middle of the working section, with a cut-out section on one wall to provide access. An acrylic panel is used to seal the tunnel, and is flush with the internal side of the wall to minimise the disturbance to the flow. Water is pumped from the plenum chamber below the working section up to the settling chamber through a return pipe, such that the flow through the working section is downwards. Seeding of the flow and the filtration of the tunnel are performed using an auxiliary circuit connected to the return pipe, which is isolated from the tunnel during experiments.
@@
Figure 1: (a) Vertical water tunnel facility and (b) experimental field of view (reproduced from [40]).
Figure 2: Aerofoil crossbeam and sting used for mounting the sphere (reproduced from [40]).
@@ 2.1.2 Sphere Design
Figure 3: Cut-away view of 3D-printed sphere models for the (a) smooth reference sphere and (c) porous spheres with and without superhydrophobic surface treatment (reproduced from [40]).
@@ 2.2 Experimental Method
The freestream velocity of the tunnel was set to $U_{\infty}=200$ mm/s for all spheres, resulting in a Reynolds number of $Re_{D}=7,780$, with a corresponding background turbulence of 2% [40]. The mean convective timescale arising from the sphere diameter and freestream velocity is $t_{c}=0.2$[40]. The flow was seeded with glass spheres of 11µm diameter, and the single-exposed particle images were illuminated by a pair of 400 mJ Spectra Physics Nd:YAG lasers firing 7 ns pulses and recorded using a PCO Panda 26 DS camera. The camera consists of a 5120 $\times$ 5120 array of 2.5 µm $\times$ 2.5 µm pixels, which was used to capture a 190 mm $\times$ 190 mm field of view beginning directly below the sphere and aligned to its centre, as shown in figure 1b, with a resulting spatial resolution of 37 µm/px. Pairs of single-exposed images were captured with a separation of $\Delta t=1$ ms, and image pairs were acquired at a rate of 1 Hz, resulting in a separation between instantaneous velocity fields of $5t_{c}$. The acquisition of $N_{T}>10,000$ images for each sphere was synchronised with the laser pulses using a BeagleBone Black (BBB) pulse generator [41].
[…]
Figure 4: Image of the plastron formed on the superhydrophobic sphere, taken with a non-PIV camera (reproduced from [40]).
@@
Freestream velocity $U_{\infty}$ 200 mm/s -
Sphere diameter $D$ 40 mm -
Reynolds number $Re_{D}$ 7,780 -
@@
Interrogation window - 64 px $\times$ 32 px 0.06DD $\times$ 0.03DD
Grid spacing - 32 px $\times$ 16 px 0.03DD $\times$ 0.015DD
Field of view - 190 mm $\times$ 190 mm 4.7DD $\times$ 4.7DD
Number of 2C-2D velocity fields $N_{T}$ >>10,000
The uncertainty of the 2C-2D-MCCDPIV algorithm used is 0.06 px [42] at the 95% confidence level, which corresponds to a single-sample uncertainty of 2.2 mm/s in the instantaneous velocity $u$. The mean of 10,000 instantaneous velocity vectors, $\overline{u}$, therefore has an uncertainty of 0.04 mm/s, accounting for the streamwise Reynolds stress [43]. As the instantaneous velocity fluctuations are given by
@@
and is equal to 1.2% of the freestream velocity at the 95% confidence level.
@@ 2.3 Dynamic Mode Decomposition
Modal decomposition is commonly used to determine the time information in patterns in turbulent fluctuations, which may indicate coherent structures and their timescale in the flow. The basis of modal decomposition is the singular value decomposition (SVD), which is also the basis of proper orthogonal decomposition (POD[44, 45]. The SVD of the $m\times n$ matrix $\boldsymbol{X}$, where $\boldsymbol{X}_{m,n}$ represents a scalar field at time $t_{m}$ and spatial coordinate $\boldsymbol{x}_{n}$, is given by
@@
As each row of both $\boldsymbol{X}_{x}$ and $\boldsymbol{X}_{x+\Delta x}$ represents an independent realisation of flow, as in the dual-PIV application by Chaugule et al. [49] to a turbulent jet. However, in the current work the separation between the vector fields is in space rather than time. As this approach is spatial rather than temporal, it does not provide a model for the temporal evolution of the flow. Instead these modes represent the spatial evolution of the flow, with the eigenvalues representing the spatial decay and wavelength of the modes.
The matrix $\boldsymbol{X}_{x}$ is defined using the instantaneous velocity fluctuations at all streamwise locations except for the furthest downstream, where each row represents an instant in time and each column represents a position in space. In order for both velocity components to be included simultaneously, the matrix $\boldsymbol{X}_{x}$ is defined as
@@
where $t_{m}$ is the time of the $m^{th}$ velocity field, ${(x,y)}_{n}$ is the $n^{th}$ spatial coordinate, and $N$ denotes the total number of spatial coordinates excluding those at the furthest downstream location. The matrix $\boldsymbol{X}_{x+\Delta x}$ is thus defined by replacing $x$ with $x+\Delta x$ in equation (10).
The reduced SVD of $\boldsymbol{X}_{x}$ is given by
@@
where $r$ is the rank of the reduction. The matrix $\boldsymbol{\widetilde{A}}$ is computed using equation (8), with $\boldsymbol{X}_{x}$ and $\boldsymbol{X}_{x+\Delta x}$ used in place of $\boldsymbol{X}_{t}$ and $\boldsymbol{X}_{t+\Delta t}$, respectively.
The eigendecomposition of $\boldsymbol{\widetilde{A}}$
@@
yields eigenvalues $\widetilde{\mu}$ and eigenvectors $\boldsymbol{\Phi}$, from which the $k^{th}$ DMD mode of $\boldsymbol{\widetilde{A}}$ is given by
@@
The eigenvalues of $\boldsymbol{A}$, $\mu$, are related to those of $\boldsymbol{\widetilde{A}}$ by
@@
The amplitudes of each DMD mode for each realisation are given by [49]
$b_{k}(t)=\boldsymbol{X}_{x+\Delta x}\boldsymbol{V}_{r}^{T}\boldsymbol{\Sigma}_{r}^{-1}\boldsymbol{\Phi}_{r}\widetilde{\mu}_{k}^{-1}.$ (15)
The scale-invariant energy contribution of each mode is given by
[…]
where
$\ldots||_{2}=\sqrt{{(\ldots)}^{*}(\ldots)}$ (17)
is the Euclidean norm, with denoting the conjugate transpose. The energy share of each mode is
$s_{k}=\frac{E_{k}}{\sum_{k=1}^{r}E_{k}}.$ (18)
@@
where $|\widetilde{\mu}_{k}|$ and $\theta_{\widetilde{\mu}_{k}}$ are the magnitude and angle of $\widetilde{\mu}_{k}$ in the complex plane, respectively. The wavenumber defined in equation (20) is in radians per metre and is divided by $2\pi$ to be converted to m-1. Therefore, the wavelength of the $k^{th}$ DMD mode in the streamwise direction, $\lambda^{x}_{k}$ is
@@
2.4 Application of Dynamic Mode Decomposition to the Flow Over Spheres
In order to capture the dynamics of the shear layer on either side of the sphere, the DMD was performed in the domains $y/D\in[0,1]$ and $y/D\in[-1,0]$ independently. This centres each domain on the corresponding side of the sphere, and maintaining the independence of the analysis of each side of the sphere means that the assumption of axisymmetry is not required. This is particularly important in the case of the PRS and SHS spheres, where the addition of the pores and superhydrophobic surface treatment may result in asymmetries.
As the flow in the wake of the sphere is highly three-dimensional, rank reduction of the SVD was necessary to limit the DMD to the most dominant modes in the measurement plane and avoid highly three-dimensional and small-scale modes that are not well-resolved by the measurement plane. Thus, the rank of the SVD was reduced to the POD modes which contributed a cumulative turbulent kinetic energy (TKE) of 25% to 50% in 5% increments. The energy share of the DMD modes at each rank was determined from their amplitudes using equations (16) to (18). The proportion of DMD modes in both domains with an energy share of greater than 5% for each case is shown for each TKE in figure 5.
Figure 5: Proportion of DMD modes with an energy share of 5% or more for each cumulative TKE for the (a) REF, (b) PRS and (c) SHS spheres. Black dotted lines indicate the chosen TKE for each sphere.
The proportion of modes with an energy share of 5% drops off sharply for a cumulative TKE of 25% for the REF and PRS spheres and 30% for the SHS sphere, shown in figure 5 for each sphere. Thus, the corresponding rank chosen for the DMD of the SVD for all three spheres was 12. The resulting rank reduction ratios were 0.21, 0.20 and 0.15 for the REF, PRS and SHS spheres, respectively. These parameters are summarised in table 2.
Table 2: Table of DMD parameters for each sphere.
Dataset TKE Ratio Rank RRR
REF 0.25 12 0.21
PRS 0.25 12 0.20
SHS 0.30 12 0.15
The DMD results for the domains above and below the centreline were related to one another by matching eigenvalues using the Hungarian method [50]. Once the corresponding modes in each domain were identified, the cross-domain complex correlation coefficient
[…]
where + and - denote the mode shapes above and below the centreline, respectively, was calculated. The angle $\angle\mathcal{R}_{+,-}^{\mathbb{C}}$ and magnitude $|\mathcal{R}_{+,-}^{\mathbb{C}}|$ of the cross-domain complex correlation coefficient gives the phase-offset and corresponding cross-correlation coefficient, respectively. The DMD mode shapes, eigenvalues and energy shares were then determined by the average of the corresponding modes in each plane.
The energy share was used to rank the modes for each sphere, and the resulting eigenvalues and energy shares are shown in figure 6. The out-of-plane vorticity
$\omega=\frac{\partial v}{\partial x}-\frac{\partial u}{\partial y},$ (23)
of each mode was calculated from its streamwise $u$ and transverse $v$ components to provide an additional visualisation of the structure of each mode. In order to find comparable modes between the spheres, the complex correlation coefficient between the $i^{th}$ and $j^{th}$ DMD modes of each sphere
[…]
was determined, where $\psi_{i}$ and $\psi_{j}$ include both the streamwise and transverse fluctuations of the corresponding modes. Modes with a wavelength longer than twice the measurement domain were moved to the end of the mode ranking, as these modes may not represent dynamics resolvable within the current measurement domain.